x+2=9
-x+3y=11
I need to know the ordered pair solution. I keep obtaining inaccurate answers.
Please help!

Answer by nerdybill(7384)

I think you have a typo in your submission.
Instead of:
x +2=9
-x+3y=11
.
You probably meant:
x +2y=9
-x+3y=11
.
Using the "addition" method we combine both equations together:
x +2y=9
-x+3y=11
------------
5y=20
y=4
.
We substitute the above into:
x +2y=9
x +2(4)=9
x +8=9
x = 1
.
Solution:(x,y) = (1,4)